# Automated Modeling Method for Pathloss Model Discovery

### v2

@@ VI-A Learning the ABG Model from Synthetic Data
(a) Train loss (the smaller, the better).
(b) Test loss (the smaller, the better).
Figure 9: Parameter search for the KAN [4,4,1] architecture for approximating $PL^{CI}$.
VI-B Learning the CI Model from Synthetic Data
Figure 8 shows the KAN architecture utilized for $PL^{CI}$ modeling, featuring a three-layer design [4, 4, 1]. The first layer consists of four input nodes corresponding to the parameters $f,n,d,\chi$, followed by a hidden layer with four nodes, and a single-node output layer representing the predicted function value. As before, to optimize the hyperparameters for KAN, we investigated various grid sizes, steps, and lambda values, as shown in Figures 9 (a and b). The 3D scatter plots illustrate how variations in the parameters affect both the train and test loss.
From the learned graph in Figure 8, it can be seen that the final expression is the summation of three terms, where the second one has a lower weight. The functions associated with the first two edges of the last layer apply nonlinear transformations, whereas the third edge performs a quasi-linear transformation. From the coloring of the edges and summations in the hidden layer, it can be observed that, in the first summation node, a quasi-linear transformation of $n$ and a non-linear transformation of $d$ are combined, both with medium weighting. The second summation node aggregates the non-linear transformation of $n$ and $d$ with low weights. Finally, the third summation node aggregates all four inputs with high to medium weights.
The expressions generated by KAN and DSRs for the $PL^{CI}$ demonstrate the validity of those modeled by KAN [4,4,1] Auto-symbolic and DSR-RSPG, as in both cases the PL is appropriately modeled as a function of $f$ and $d$. We note that, same as in the case of the $PL^{ABG}$, the expression produced by the KAN [4,4,1] Auto-symbolic is quite complex. The expression generated by DSR-PQT, although it includes both $f$ and $d$, models the frequency term using $\cos(\log_{10}(f))$. This implies a periodic contribution of $f$ to PL, which contradicts the fundamental characteristics of signal propagation, as PL increases with increasing $f$. Furthermore, the expression obtained by DSR-VPG excludes both $f$ and $d$. These results imply the invalidity of the DSR-PQT and DSR-VPG expressions for the PL modeling.
@@ VI-C Learning an Indoor Model from Empirical Data
The performance comparison of various methods for modeling PL in an indoor environment is presented in Table VII. The $PL^{MWF}$ serves as the baseline, yielding the highest errors (MAE = 9.01, MSE = 121.54, MAPE = 8.98) and $R^{2}$ value of 0.56, indicating a moderate predictive performance. $PL^{EI}$, based on Eq. (4) from [35], slightly improves accuracy, with MAE = 7.64, MSE = 111.67, MAPE = 7.52, and $R^{2}$ = 0.49. However, all automated methods outperform both the baseline and the empirical model. KAN [4,1] significantly improves accuracy (MAE = 6.11, MSE = 59.87, MAPE = 6.22) and a substantially higher $R^{2}$ value of 0.76, as shown in Figure 10, where pred stands for the predicted PL values and true stands for the actual measurements from [35]. The KAN [4,1] Auto-Symbolic performs similarly, with slightly higher errors but maintaining the same $R^{2}$ value. Among DSR methods, DSR-RSPG achieves the overall superior performance (MAE = 5.95, MSE = 55.73, MAPE = 5.96, $R^{2}$ = 0.71), though its $R^{2}$ remains inferior to that of the KAN models. DSR-PQT and DSR-VPG demonstrate comparatively weaker performance, with higher errors and lower values of $R^{2}$ (0.65 and 0.57, respectively). Overall, the results highlight that both KANs and DSR models provide substantial improvements over the baseline model and the indoor empirical model.
Figure 11 illustrates the architecture of the KAN [4,1] used for indoor PL prediction. The architecture has two layers; the first layer consists of four input nodes corresponding to the parameters $n_{w},n_{f},d,f$, and a single-node output layer representing the predicted value. As illustrated in the figure, the final expression is a sum of four terms, with the fourth term contributing less due to its smaller weight. Notably, the function learned on the first edge applies a non-linear transformation, whereas the remaining three edges implement quasi-linear transformations.
@@
Methods Metrics MAE MSE MAPE R2 Expression
$\boldsymbol{PL^{MWF}}$ 9.01 121.54 8.98 0.56 $10n\log_{10}\left(d\right)+PL_{0}+\mathrm{WAF}+\mathrm{FAF}$
$\boldsymbol{PL^{EI}}$ 7.64 111.67 7.52 0.49 $10n\log_{10}(d)+PL_{0}+n_{w}L_{w}+n_{f}^{\left(\frac{n_{f}+2}{n_{f}+1}-b\right)}L_{f}$
KAN [4,1] 6.11 59.87 6.22 0.76 learned tree and splines as in Figure 11.
KAN [4,1] Auto-symbolic 6.27 61.62 6.38 0.76 $41.516\log_{10}\left(0.383d+4.03\right)-1.328\cos\left(5.036f-0.448\right)$ $+4.912\sin\left(8.226n_{w}-2.854\right)+52.026\log_{10}\left(2.408n_{f}+8.63\right)-11.236$
DSR-RSPG 5.95 55.73 5.96 0.71 $\frac{d}{2}+3log_{10}(f)+5n_{w}+70$
DSR-PQT 7.09 80.99 7.19 0.65 $\frac{d}{2}+4n_{w}+80$
DSR-VPG 7.38 84.11 7.29 0.57 $n_{f}^{2}+4n_{f}+n_{f}n_{w}+80$
@@ VI-D Learning an Outdoor Model from Empirical Data
Table VIII summarizes the performance of different methods for PL prediction in an outdoor environment. The $PL^{FS}$, yields high error values (MAE = 34.08, MSE = 1242.99, MAPE = 31.43) and a low $R^{2}$ value of 0.30, indicating poor predictability. The $PL^{EO}$, represented by Eq. (5) from [35], substantially improves accuracy (MAE = 18.09, MSE = 398.88, MAPE = 17.82), but retains the same $R^{2}$ value of 0.30. In contrast, the KAN model [3,1] achieves significantly better performance, with the lowest errors (MAE = 4.59, MSE = 38.59, MAPE = 4.36) and a higher $R^{2}$ value of 0.71, as demonstrated in Figure 12 showing the comparison between predicted and measured PL values. The KAN [3,1] Auto-symbolic yields poor performance (MAE = 9.50, $R^{2}$ = 0.01), showing a higher error and a lower fit compared to the original KAN structure. Among the DSR models, DSR-RSPG performs the best (MAE = 6.62, MSE = 72.09, MAPE = 6.31); however, its $R^{2}$ value (0.29) remains low. DSR-PQT and DSR-VPG show slightly higher errors and comparable $R^{2}$ values. Overall, the KAN [3,1] model outperforms both traditional and DSR methods, offering a more accurate method for outdoor PL prediction.
Figure 13 illustrates the KAN architecture evaluated for the outdoor model, consisting of a two-layer design [3, 1]. The first layer includes three input nodes corresponding to the parameters $h_{\text{ED}},d,f$, followed by a single output node representing the predicted PL value. The figure shows that the final expression is a sum of three terms, with the first and third terms having smaller weights. It can be observed that the functions learned on the second and third edges apply non-linear transformations, whereas the first implements a quasi-linear transformation.
@@
Methods Metrics MAE MSE MAPE R2 Expression
$\boldsymbol{PL^{FS}}$ 34.08 1242.99 31.43 0.30 $32.4420\log_{10}(d)+20\log_{10}(f)+$ 32.44
$\boldsymbol{PL^{EO}}$ 18.09 398.88 17.82 0.30 $10n\log_{10}(d)+PL_{0}+L_{h}\log_{10}(h_{\text{ED}})+X_{\sigma}$
KAN [3,1] 4.59 38.59 4.36 0.71 learned tree and splines as in Figure 13
KAN [3,1] Auto-symbolic 9.50 133.42 9.34 0.01 $105.740.99\cos{\left(8.71f-3.47\right)}-2.34\operatorname{log}_{10}{\left(7.68-2.16h_{\text{ED}}\right)}+$ 105.74
DSR-RSPG 6.62 72.09 6.31 0.29 $h_{\text{ED}}d+\log_{10}(f)+80$
DSR-PQT 7.25 80.45 6.78 0.29 $\log_{10}(\frac{d}{h_{\text{ED}}})+\frac{d}{10}+90$
DSR-VPG 8.18 97.90 7.46 0.28 $\frac{d+sin(f)}{10}+90$
The only valid expression for the outdoor environment is obtained from DSR-RSPG, as it includes both $f$ and $d$. The KAN [3,1] Auto-symbolic expression excludes $d$ and incorporates a cosine function of $f$. The expression of DSR-PQT excludes $f$. Finally, although the DSR-VPG expression includes both $f$ and $d$, it defines $f$ using a sine function, introducing a periodicity that misrepresents the actual relationship between $f$ and PL.
@@ VI-E Discussion
In general, our findings indicate that KANs provide highly accurate solutions with higher interpretability compared to typical activation maps of DNNs, but lower than some of the considered DSR methods. On the other hand, DSR methods prioritize interpretability over predictive accuracy. The choice between these methods should be determined by the intended objective, whether to maximize predictive accuracy or to derive interpretable expressions that offer clear analytical insight.
@@ VII Conclusion
In future work, our goal is to focus on enhancing the symbolic mapping mechanisms within KANs for improved interpretability. Furthermore, extending both the KANs and DSR methods to cover diverse scenarios will be essential to validate their robustness and adaptability. These advances have the potential to contribute significantly to the development of intelligent, adaptive modeling frameworks in the field of wireless communication systems.

### v3

@@ VI-B Learning the CI Model from Synthetic Data
TABLE VI: Performance comparison on CI model.
Methods Metrics MAE MSE MAPE R2 Expression Interpretability Validity
KAN [4,4,1]* 0.05 0.01 0.04 1 Tree and splines in Figure 7 Medium ?
KAN [4,4,1] Auto-symbolic* 4.56 36.26 3.05 0.96 $-280.5\operatorname{log}_{10}{\left(8.7-1.6n\right)}+43.8\operatorname{log}_{10}{\left(1.1d+1.9\right)}+19.9\operatorname{log}_{10}{\left(5.2f+9.3\right)}+50.8\operatorname{log}_{10}{\left(0.8\chi_{\sigma}+7.5\right)}+352.2$ High
DSR-RSPG 11.49 261.46 10.21 0.80 $23n+\frac{d}{10}+log_{10}(f)+40$ High
DSR-PQT 13.60 223.05 10.58 0.84 $23n+\frac{d}{10}+(\chi_{\sigma}+n)cos(log_{10}(f))+40$ High
DSR-VPG 16.90 570.92 15.24 0.67 $21n+2\chi_{\sigma}+60$ High
* Input parameters were normalized by dividing each value by the maximum value of its corresponding parameter
Interpretability
Figure 7 shows the KAN architecture utilized for $PL^{CI}$ modeling, featuring a three-layer design [4, 4, 1]. The first layer consists of four input nodes corresponding to the parameters $f,n,d,\chi_{\sigma}$, followed by a hidden layer with four nodes, and a single-node output layer representing the predicted function value.
From the learned graph in Figure 7, it can be seen that the final expression is the sum of three terms, where the second has a lower weight. The rightmost node is disconnected because the model has learned that the connecting edges are not important. The functions associated with the first two edges of the last layer apply nonlinear transformations, whereas the third edge performs a quasi-linear transformation. From the coloring of the edges and summations in the hidden layer, it can be observed that, in the first summation node, a quasi-linear transformation of $n$ and a nonlinear transformation of $d$ are combined, both with medium weighting. The second summation node aggregates the nonlinear transformation of $n$ and $d$ with low weights. Finally, the third summation node aggregates all four inputs with high to medium weights.
Validity
The expressions generated by KAN and DSRs for the $PL^{CI}$ demonstrate the validity of those modeled by KAN [4,4,1] Auto-symbolic and DSR-RSPG, as in both cases the PL is appropriately modeled as a function of $f$ and $d$. We note that, same as in the case of the $PL^{ABG}$, the expression produced by the KAN [4,4,1] Auto-symbolic is quite complex, while the composition graph requires additional investigation. The expression generated by DSR-PQT, although it includes both $f$ and $d$, models the frequency term using $\cos(\log_{10}(f))$. This implies a periodic contribution of $f$ to PL, which contradicts the fundamental characteristics of signal propagation, as PL increases with increasing $f$. Furthermore, the expression obtained by DSR-VPG excludes both $f$ and $d$. These results imply the invalidity of the DSR-PQT and DSR-VPG expressions for the PL modeling.
@@ VI-C Learning an Indoor Model from Empirical Data
The performance comparison of various methods for modeling PL in an indoor environment is presented in Table VII. The $PL^{MWF}$ serves as the baseline, yielding the highest errors (MAE = 9.01, MSE = 121.54, MAPE = 8.98) and $R^{2}$ value of 0.56, indicating a moderate predictive performance. $PL^{EI}$, based on Eq. (4) from [ElChall], slightly improves accuracy, with MAE = 7.64, MSE = 111.67, MAPE = 7.52, and $R^{2}$ = 0.49. However, all automated methods outperform both the baseline and the empirical model. KAN [4,1] significantly improves accuracy (MAE = 6.11, MSE = 59.87, MAPE = 6.22) and a substantially higher $R^{2}$ value of 0.76, as shown in Figure 8, where pred stands for the predicted PL values and true stands for the actual measurements from [ElChall]. The KAN [4,1] Auto-Symbolic performs similarly, with slightly higher errors but maintaining the same $R^{2}$ value. Among DSR methods, DSR-RSPG achieves the overall superior performance (MAE = 5.95, MSE = 55.73, MAPE = 5.96, $R^{2}$ = 0.71), though its $R^{2}$ remains inferior to that of the KAN models. DSR-PQT and DSR-VPG demonstrate comparatively weaker performance, with higher errors and lower values of $R^{2}$ (0.65 and 0.57, respectively). Overall, the results highlight that both KANs and DSR models provide substantial improvements over the baseline model and the indoor empirical model.
Interpretability
Figure 9 illustrates the architecture of the KAN [4,1] used for indoor PL prediction. The architecture has two layers; the first layer consists of four input nodes corresponding to the parameters $n_{w},n_{f},d,f$, and a single-node output layer representing the predicted value. As illustrated in the figure, the final expression is a sum of four terms, with the fourth term contributing less due to its smaller weight. Notably, the function learned on the first edge applies a nonlinear transformation, whereas the remaining three edges implement quasi-linear transformations.
@@
Methods Metrics MAE MSE MAPE R2 Expression Interpretability Validity
$\boldsymbol{PL^{MWF}}$ Eq. 6* 9.01 121.54 8.98 0.56 $10n\log_{10}\left(d\right)+PL_{0}+\mathrm{WAF}+\mathrm{FAF}$ High
$\boldsymbol{PL^{EI}}$ Eq. 4* 7.64 111.67 7.52 0.49 $10n\log_{10}(d)+PL_{0}+n_{w}L_{w}+n_{f}^{\left(\frac{n_{f}+2}{n_{f}+1}-b\right)}L_{f}$ High
KAN [4,1] 6.11 59.87 6.22 0.76 learned tree and splines as in Figure 9. Medium ?
KAN [4,1] Auto-symbolic 6.27 61.62 6.38 0.76 $41.5\log_{10}\left(0.4d+4.03\right)-1.3\cos\left(5.04f-0.5\right)$ $4.9\sin\left(8.2n_{w}-2.9\right)+52.03\log_{10}\left(2.4n_{f}+8.6\right)-11.2$ High
DSR-RSPG 5.95 55.73 5.96 0.71 $\frac{d}{2}+3log_{10}(f)+5n_{w}+70$ High
DSR-PQT 7.09 80.99 7.19 0.65 $\frac{d}{2}+4n_{w}+80$ High
DSR-VPG 7.38 84.11 7.29 0.57 $n_{f}^{2}+4n_{f}+n_{f}n_{w}+80$ High
* Expressions developed with traditional methods used as baselines.
Validity
@@ VI-D Learning an Outdoor Model from Empirical Data
Table VIII summarizes the performance of different methods for PL prediction in an outdoor environment. The $PL^{FS}$, yields high error values (MAE = 34.08, MSE = 1242.99, MAPE = 31.43) and a low $R^{2}$ value of 0.30, indicating poor predictability. The $PL^{EO}$, represented by Eq. (5) from [ElChall], substantially improves accuracy (MAE = 18.09, MSE = 398.88, MAPE = 17.82), but retains the same $R^{2}$ value of 0.30. In contrast, the KAN model [3,1] achieves significantly better performance, with the lowest errors (MAE = 4.59, MSE = 38.59, MAPE = 4.36) and a higher $R^{2}$ value of 0.71, as demonstrated in Figure 10 showing the comparison between predicted and measured PL values. The KAN [3,1] Auto-symbolic yields poor performance (MAE = 9.50, $R^{2}$ = 0.01), showing a higher error and a lower fit compared to the original KAN structure. Furthermore, the very low $R^{2}$ score shows low interdependence between the input and output variables, meaning that the function to spline matching process resulted in an inaccurate match, leading to significant degradation relative to the KAN-only network.
Among the DSR models, DSR-RSPG performs the best (MAE = 6.62, MSE = 72.09, MAPE = 6.31); however, its $R^{2}$ value (0.29) remains low. DSR-PQT and DSR-VPG show slightly higher errors and comparable $R^{2}$ values. Overall, the KAN [3,1] model outperforms both traditional and DSR methods, offering a more accurate method for outdoor PL prediction.
Interpretability
Figure 11 illustrates the KAN architecture evaluated for the outdoor model, consisting of a two-layer design [3, 1]. The first layer includes three input nodes corresponding to the parameters $h_{\text{ED}},d,f$, followed by a single output node representing the predicted PL value. The figure shows that the final expression is a sum of three terms, with the first and third terms having smaller weights. It can be observed that the functions learned on the second and third edges apply nonlinear transformations, whereas the first implements a quasi-linear transformation.
@@
Methods Metrics MAE MSE MAPE R2 Expression Interpretability Validity
$\boldsymbol{PL^{FS}}$ Eq. 7* 34.08 1242.99 31.43 0.30 $20\log_{10}(d)+20\log_{10}(f)+32.44$ High
$\boldsymbol{PL^{EO}}$ Eq. 5* 18.09 398.88 17.82 0.30 $10n\log_{10}(d)+PL_{0}+L_{h}\log_{10}(h_{\text{ED}})+X_{\sigma}$ High
KAN [3,1] 4.59 38.59 4.36 0.71 learned tree and splines as in Figure 11 Medium ?
KAN [3,1] Auto-symbolic 9.50 133.42 9.34 0.01 $\cos{\left(8.7f-3.5\right)}-2.3\operatorname{log}_{10}{\left(7.7-2.2h_{\text{ED}}\right)}+105.8$ High
DSR-RSPG 6.62 72.09 6.31 0.29 $h_{\text{ED}}d+\log_{10}(f)+80$ High
DSR-PQT 7.25 80.45 6.78 0.29 $\log_{10}(\frac{d}{h_{\text{ED}}})+\frac{d}{10}+90$ High
DSR-VPG 8.18 97.90 7.46 0.28 $\frac{d+sin(f)}{10}+90$ High
* Expressions developed with traditional methods used as baselines.
Validity
The only valid expression for the outdoor environment is obtained from DSR-RSPG, as it includes both $f$ and $d$. The KAN [3,1] Auto-symbolic expression excludes $d$ and incorporates a cosine function of $f$, which implies periodicity that is not the case with pathloss models. The lack of validity introduced by this cosine component is also confirmed by the very low $R^{2}$ score. The expression of DSR-PQT excludes $f$. Finally, although the DSR-VPG expression includes both $f$ and $d$, it defines $f$ using a sine function, introducing a periodicity that misrepresents the actual relationship between $f$ and PL.
@@ VI-E Discussion
Overall, our findings indicate that KANs deliver highly accurate results with superior first-level interpretability(spline-based structures) compared to typical DNN activation maps. However, when it comes to the second level of interpretability, where actual functions are extracted using the Auto-symbolic, KANs are outperformed by certain DSR methods. On the other hand, DSR methods prioritize interpretability over predictive accuracy. The choice between these methods should be determined by the intended objective, whether to maximize predictive accuracy or to derive interpretable expressions that offer clear analytical insight.
@@ VII Conclusion
This work also revealed several challenges to be addressed in the future. More specifically, utilizing a clear form of path loss, the non-logarithmic form, may enable more precise modeling. The performance-interpretability tradeoffs with KANs could likely be improved by enhanced symbolic mapping. Further methodological tools to assess the validity of the composition graph learned by the KAN may ease the adoption of the proposed method. Finally, extending both the KANs and DSR methods to cover diverse scenarios will be essential to validate their robustness and adaptability. These advances have the potential to contribute significantly to the development of intelligent, adaptive modeling frameworks in the field of wireless communication systems.
Acknowledgements
This work was supported by the Slovenian Research Agency (P2-0016) and the European Commission NANCY project (No. 101096456).